FA 1 -PHUONG CASE STUDY: TOPIC 8 Following his decision not to go ahead with either of the expansion plans you discussed with him before, Phuong has clarified with you that financial position of the business has remained unchanged and is outlined below Assets -112,623 - 51,787 (this consists of the 47,024 loan from the bank and the remainder being existing liabilities held by the business) ー60,836 Liabilities Owners Equity Phuong has a cousin named Hoang who is very keen to be involved in the business and thinks the idea of going into partnership would be great. From Hoangs point of view he thinks he and Phuong should be equal partners with a half share each in the business. Hoang has told Phuong that he will contribute the following assets and this should allow him to obtain his half share. All assets are listed at their fair value Cash - $20,000, Motor Vehicle - $18,000, Computer Equipment - $9,000. Required Determine the impact of allowing Hoang to have a half share of the business by contributing these assets and show the journal entries if this was to take place. Also comment on whether you think this is a reasonable offer for Hoang to accept and why you have come to this conclusion 2.

ph 2 36 2 a ort h (7346A 173 4 huong felt like he had little choice but to work with Hoang as he is family and that while he had some concerns the extra injection of capital would help. As Phuong suspected it was h ard work getting Hoang to make a real contribution, and even turning up was an issue at times but despite this the business managed to generate a reasonable profit of $23,000. Phuong has asked Before starting the new partnership Phuong agreed that being family they would distribute any know anything about how the business was run and didnt really want to learn, preferring just to modest Annual Salary of $15,000. Hoang agreed to this on the condition that to reward him for you to help him work out how this profit should be distributed amongst the partners profits/losses on a 50:50 basis. However, Phuong quite cleverly made the point that Hoang didnt give his expert advice and that he would be willing to run the day to day operations for a contributing his hard earned money to the business he would get a return of 10% his capital balance but Phuong would only receive 500 as he was already getting a was all agreed to and put in the partnership agreement interest on salary. This Required 3. Calculate the profit distribution after the first year and show how the Equity section of he balance sheet would look after this distribution. Caton 23,0 jan
Answer #1

Total capital of the business before admission of new partner = $60836.

Required capital after admission of new partner = 60836/50% = $121672

Investment required from new partner = 121672-60836 = $60836.

Investment made by Hoang = 20000+18000+9000 = $47000

This is a reasonable offer Hoang to accept since the fair value of assets he is contributing is coming out to be less than what he was required to contribute

Journal entry will be made considering Goodwill since there is difference between paid in capital and required capital .

From above we came to know that partnership valuation is $121672. However total paid in capital = $60836+$49000 = $109836.

Goodwill = 121672-109836 = $11836

Since new partner is gaining the entire Goodwill will belong to him

Journal entry will be as follows :-

General journal debit Credit

Cash

Motor vehicle

Computer equipment

Goodwill

Haong capital

(To admission of haong)

20000

18000

9000

11836

58836

Prifit will be distributed after salary and interest.